The time it takes to preform a task has a continuous uniform distribution between 43 min and 57 min. What is the the probability it takes between 51.4 and 54.5 min. Round to 4 decimal places. P(51.4 < X < 54.5) can you explain this too please
Answer #1

Ans:

Uniform distribution with a=43 min and b=57 min

P(X<=x)=(x-a)/(b-a)

probability that it takes between 51.4 and 54.5 min=P(51.4<X<54.5)

=P(X<=54.5)-P(X<=51.4)

=(54.5-43)/(57-43)-(51.4-43)/(57-43)

=(54.5-51.4)/(57-43)

=0.2214